a bakery sold 300 cupcakes in 12 days then sold 450 cupcakes in 18 days. bakery b sold 120 cupcakes in 5 days, then sold 200 cupcakes in 8 days. which bakery had the greatest rate of increase? The rate of increase is found using slope.you are looking for the increase in cupcakes per day.
Answer by stanbon(75887)

bakery A sold 300 cupcakes in 12 days then sold 450 cupcakes in 18 days.
rate = 450/18 = 25 (cakes/day)
-------------------------------------------
bakery B sold 120 cupcakes in 5 days, then sold 200 cupcakes in 8 days.
rate = (120+200)/(5+8) = 320/13 = 24.61 (cakes/day)
---------------------------------------------
which bakery had the greatest rate of increase? The rate of increase is found using slope.you are looking for the increase in cupcakes per day.
----
Ans: A's rate is greater
--------------
Cheers,
Stan H.
-------------